1.The half-life of Iodine-131 is 8 days. Your aunt has an overactive thyroid, so her doctor gives her a pill containing 1 gram of I-131. 32 days later, how many grams remain in her body?
2.A granite counter top comes from a rock formation that is 3,060,000 years old. 50% of the original zirconium-93 is still in the rock. What is the half-life of zirconium-93?
3.What percent of Radium-226 remains after 6488 years have passed?(halflife:1622 years)
4.A block of metal once contained 2000 atoms of lead-202. After some time passed, it now holds 250 atoms. How many half lives have gone by?
1
Expert's answer
2020-03-30T10:59:15-0400

1. x = 32 / 8 = 4

m(I) = 1 / 24 = 1 / 16 = 0.0625 (g)

2. 50% = 1 / 21 "\implies" x = 1 "\implies" T1/2 = 3,060,000 (years)

3. x = 6488 / 1622 = 4

"\omega" = 1 / 24 = 0.0625 = 6.25 %

4. x = log2(2000 / 250) = log2(8) = 3